Christine and Peter earn a combined annual income of $290,500. If Christine earns $20,500 per year more than Peter, what are their individual annual salaries?
Answer by checkley79(3341)

LET X BE PETER'S SALARY
LET X+20,500 BE CHRISTINE'S SALARY
X+X+20.500=290,500
2X=290,500-20,500
2X=270,000
X=270,000/2
X=$135,000 IS PETER'S SALARY
135,000+20,500=$155,500 IS CATHERINE'S SALARY.
PROOF:
135,000+155,500=$290,500